OpenStudy (anonymous):

solve this system with elimination -x+5y=13 x-y=15

OpenStudy (mertsj):

If you add the two equations the x terms will drop out since they are opposites.

OpenStudy (nikato):

If u add them both up, u'll get 4y=28 So what is y?

OpenStudy (anonymous):

-x+5y=13 + x-y=15 add them together, and the x's cancel. 4y=28. Divide by 4 on both sides, y=7. Plug 7 into one of the equations. x-(7)=15 +7 +7 x=22 The solution set is (22,7)